Cutter Enterprises purchased equipment for $48,000 on January 1, 2018. The equipment is expected to have a five-year life and a residual value of $6,600. Using the double-declining balance method, depreciation for 2018 and the book value at December 31, 2018, would be:
Answer #1

Calculation of depreciation as per double declining:

Depreciation= (original cost-salvage value)/useful life

Depreciation= (48000-6600)/5= 8280

Depreciation rate= 8280/48000*100= 17.25%

Double declining depreciation rate= 17.25*2= 34.50%

Depreciation= 48000*0.345= 16560

Depreciation is 16560

Book value= 48000-16560= 31440